OpenStudy (anonymous):

121a^6x^4 square root by 4 equation is next. please help if you can reduce the index as small as possible.

OpenStudy (anonymous):

\[\sqrt[4]{121a^6x^4}\]

OpenStudy (anonymous):

\[(121 a^{6} x^{4})^{1/4} = 121^{1/4}*a^{6/4}*x^{4/4}=\sqrt{11}a^{3/2}x\]